OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6.

OpenStudy (compassionate):

Hello, do you know the first step?

OpenStudy (anonymous):

For two lines to be perpendicular, their slopes must be negative reciprocals of each other. So, for our new line, \[m = \frac{ -1 }{ 3 }\] Now, you have the slope and a point on the line, so plug that into the equation: \[y-y_1=m(x-x_1)\] and of course, simplify if needed.

OpenStudy (anonymous):

idk how @vinnv226

OpenStudy (anonymous):

Ok, we'll start by plugging in our values: \[y- -6 = (-1/3)(x-6)\]\[y+6=(-1/3)(x-6)\]Now we'll distribute our slope: \[y+6=(-x/3)+2\]Finally, subtract 6 from both sides: \[y=(-x/3)-4\]

OpenStudy (anonymous):

Choose the equation of the vertical line that passes through the point (-5, 9). @vinnv226

OpenStudy (anonymous):

A vertical line is always of the form \[x=c\]where c is some constant. Since you know the coordinates of a point (-5,9) all the values on that vertical line have to have -5 as their x coordinate. So your equation is \[x=-5\]

OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation y - 4x = -3 @vinnv226